I need some help please.
The Nero Match Company sells matchboxes that are supposed to have an average
of 40 matches per box with σ = 9 . A random sample of 94 Nero matchboxes
shows the average number of matches per box to be 43.1. At a 1% level of
significance, can you say that the average number of matches per box is more than 40?

Answer by stanbon(75887)

The Nero Match Company sells matchboxes that are supposed to have an average
of 40 matches per box with σ = 9 . A random sample of 94 Nero matchboxes
shows the average number of matches per box to be 43.1. At a 1% level of
significance, can you say that the average number of matches per box is more than 40?
------
Ho: u = 40
Ha: u > 40
-----
z(43.1) = (43.1-40)/[9/sqrt(04)] = 3.3395
--------
p-value = P(z > 3.3395) = normalcdf(3.3395,100) = 0.000419
-----
Since the p-value is less than 1%, reject Ho at the 1%
level of significance.
Conclusion: The test results do not support saying the mean number is 40.
--------------
Cheers,
Stan H.